The length of a rectangle is twice its breadth. If its perimeter is 54 cm, find its length
Answer by checkley77(12844)

P=2L+2W
L=2W
54=2*2W+2W
54=4W+2W
6W=54
W=54/6
W=9 ANS. FOR THE WIDTH.
L=2*9=18
54=2*9+2*18
54=18+36
54=54